if there are 42g of hydrogen (H) in a sample of pure methane (cH4) how many grams of carbon (C) are present
1
Expert's answer
2020-11-18T13:41:42-0500

"n(H) = \frac{42 \;g}{1 \;g/mol} = 42 \; mol"

"n(C) = \frac{1}{4}n(H) = \frac{42 \; mol}{4} = 10.5 \;mol"

"m(C) = 10.5 \; mol \times 12 \;g/mol = 126 \;g"

Answer: 126 g